2. Solve the below truss for all reactions and member forces - 10 K 1 0 10- 10- 10 20 kip 20 kip 20 kips
Answer #1

All A lol low zok Reactions cok 2ok RAY 10:45 zo SFn = o ; han = of E Miro o-20x30-20x20-20x1050 (RAY = 30k) (1) < Ey=0; Riy=At Jontcl FBC FLD < Fy=0, FD + Fec-20=0 FCD= 20-3054 Fed=-10k) (compression) FCE FAC 20k & Fu=0; FEE - FAC - FBC To FLE - 305Art Jant IF & Ry=0; FFG & FEF=0 2 + FEFTO FOR N45 4 FIH F FFn=-1052k FFG -14.142 K FEE compression) { Furo, FFH+Ffg -For to FFHI 1 FIJ Art Jout I < Fy=0; FIS + FHE +30=0 Frs = -30 + 3017 FOR I 3ok